# Likelihood-Ratio Regularized Quantile Regression: Adapting Conformal Prediction to High-Dimensional Covariate Shifts

### v1

@@ 5 Experiments
In this section, we compare our method, the LR-QR algorithm, with the following baselines: (1) Split or inductive conformal prediction Papadopoulos et al., (2002); Lei et al., (2018); (2) Weighted conformal prediction Tibshirani et al., (2019), referred to as Weighted-CP; (3) The doubly robust method proposed in Yang et al., (2024), referred to here as 2R-CP; (4) The distributionally robust optimization approach from Cauchois et al., (2024), referred to as DRO-CP.
5.1 Choosing the regularization parameter $\lambda$
Equation 6 suggests an optimal choice of the regularization parameter $\lambda$ in LR-QR. Guided by this theory, we form a grid of ten equally spaced candidate values, ranging from the theoretical $\lambda^{*}$ down to $\frac{\lambda^{*}}{10}$. We then perform three-fold cross-validation over the combined calibration and unlabeled sets (without using any labeled test data) as follows: for each candidate $\lambda$, we train LR-QR with that $\lambda$, and compute a validation measure, namely the $\ell^{2}$-norm of the gradient of the LR-QR objective on the held-out fold. We pick $\lambda$ with the smallest average validation measure across all folds.
This validation measure is motivated by our algorithmic development: the first-order conditions of LR-QR play a fundamental role in ensuring valid coverage in the test domain. While the model is trained to satisfy these conditions on the observed data, we seek to ensure this property generalizes well to unseen data. Thus, our selection criterion is based on two key observations: (1) a small gradient of the LR-QR objective implies reliable coverage, and (2) the regularization parameter $\lambda$ balances the generalization error of the two terms in LR-QR. By minimizing this measure, we select a $\lambda$ that optimally trades off these competing factors.
Finally, we re-train LR-QR on the entire calibration and unlabeled sets using this best $\lambda$, and report coverage and interval size on the held-out labeled test set. This ensures that no test labels are used during hyper-parameter tuning.
Additionally, in Appendix B, we provide deeper insight on different regimes of regularization in practice through an ablation study.
@@ 5.2 Communities and Crime
We evaluate our methods on the Communities and Crime dataset Redmond, (2002), which contains $1994$ datapoints corresponding to communities in the United States, along with socio-economic and demographic statistics. The task is to predict the per-capita violent crime rate (a real-valued response) from a 127-dimensional input.
@@
Setup. We first randomly select half of the data as a training set, and use it to fit a ridge regression model $\hat{f}(\cdot)$ as our predictor. We tune the ridge regularizer with five-fold cross-validation. We use the remaining half to design four covariate shift scenarios, determined by the frequency of a specific racial subgroup (Black, White, Hispanic, and Asian). For each of these features, we find the median value $m$ over the remaining dataset. Datapoints with feature value at most $m$ form our source set, while the rest form our target set. This creates a covariate shift between calibration and test, as the split procedure only observes the covariates, and is independent of labels. We then further split the target set into roughly equal unlabeled and labeled subsets. The unlabeled subset and the calibration data (only their covariates, not the labels) is used for estimating $r$, whereas the labeled subset of test is held out only for final evaluation. The same procedure is applied to each of the four racial subgroups, creating four distinct partitions.
@@
Conformity score and likelihood ratio estimation. The conformity score is the absolute residual $(x,y)\mapsto s(x,y)\;=\;|y-\hat{f}(x)\,|.$ Several baselines require an estimate of the likelihood ratio $r$, which we obtain by training a logistic regression model to distinguish datapoints in the unlabeled target data from those in the source data. We then use $\hat{r}=\hat{p}/\bigl{(}1-\hat{p}\bigr{)}$, where $x\mapsto\hat{p}(x)$ is the predicted probability that $x$ belongs to the target distribution.
Experimental details. We set the function class $\mathcal{H}$ to be linear maps from the feature space to $\mathbb{R}$. The regularization $\lambda$ is tuned as explained in Section 5.1. All experimental results are averaged over 1000 random splits.
@@
Results. Figure 1 depicts the performance of all methods. Notably, split conformal undercovers in two setups and overcovers in the other two. Methods relying on estimating the likelihood ratio and DRO fail to track the nominal coverage precisely, particularly in the first setup on the left. However, LR-QR is closer to the nominal level of coverage, showing a stronger adaptivity to the covariate shift.
5.3 RxRx1 data - WILDS
Our next experiment uses the RxRx1 dataset Sypetkowski et al., (2023) from the WILDS repository Koh et al., (2021), which is designed to evaluate model robustness under distribution shifts. The RxRx1 task involves classifying cell images based on 1,339 laboratory genetic treatments. These images, captured using fluorescent microscopy, originate from 51 independent experiments. Variations in execution and environmental conditions lead to systematic differences across experiments, affecting the distribution of input features (e.g., lighting, cell morphology) while the relationship between inputs and labels remains unchanged. This situation creates covariate shift where the marginal distribution of inputs shifts across domains, but the conditional distribution $\mathbb{P}_{Y|X}$ remains the same.
For this task, we use a ResNet50 model He et al., (2016) trained by the WILDS authors on 37 of the 51 experiments. Using data from the other 14 experiments, we construct 14 distinct evaluations. In each evaluation, one experiment is selected as the target dataset, and its data is evenly split into an unlabeled target set and a labeled test set. The labeled data from the other 13 experiments serves as the source dataset. We evaluate all baselines across these 14 settings.
Conformity score and likelihood ratio estimation. The nonconformity score $S(x,y)$ is the negative log-likelihood $(x,y)\mapsto-\log f_{x}(y),$ where $f_{x}(y)$ is the probability the pretrained model assigns to the image-label pair $(x,y)$. Several baselines require an estimate of $r$, which we obtain by training a logistic regression model on top of the representation layer of the pretrained model to distinguish datapoints in the unlabeled target data from those in the source data. We then use $\hat{r}=\hat{p}/\bigl{(}1-\hat{p}\bigr{)}$, where $x\mapsto\hat{p}(x)$ is the predicted probability that $x$ came from the target distribution.
Experimental details. For our proposed method, we set the function class $\mathcal{H}$ to be a linear head on top of the representation layer of the pretrained model. The $\lambda$ is tuned using the cross-validation explained in Section 5.1. Also, all experimental results are averaged over 50 random splits.
Results. Figure 2 presents the coverage and average prediction set size for all methods. To enhance visual interpretability, we display results for eight randomly selected settings out of the 14, with the full plot provided in Figure 3 in the Appendix. The x-axis shows the indices of the test experimental condition. As shown, LR-QR consistently adheres more closely to the nominal coverage value of $0.9$ compared to other baselines. The average absolute deviations of coverage from the nominal value in terms of percentage for LR-QR, 2R-CP, weighted-CP, DRO-CP, and Split conformal are $1.56,2.62,4.03,5.25$, and $5.36$, respectively.
Notably, split conformal prediction, which assumes exchangeability between calibration and test data, shows under- and overcoverage due to the shift. The coverage of weighted CP and 2R-CP is also far from the target, showing that directly estimating the likelihood ratio and conditional quantile is insufficient to correct the coverage violations in the case of high-dimensional image data. Further, the superior coverage of LR-QR is not due to inflated prediction sets, as the sizes are on par with other methods.
@@ Appendix A Additional figures
Figure 3: (Left) Coverage, (Right) Average prediction set size.
@@
Appendix B Ablation studies
@@
Appendix C Further related work
@@
Appendix D Notation and conventions
Constants are allowed to depend on dimension only through properties of the population and sample covariance matrices of the features, and the amount of linear independence of the features; see the quantities $\lambda_{\textnormal{min}}(\Sigma)$, $\lambda_{\textnormal{max}}$, $c_{\textnormal{min}}$, $c_{\textnormal{max}}$, and $c_{\textnormal{indep}}$ defined in Appendix E. In Landau notation ($o$, $O$, $\Theta$), we hide constants. We say that a sequence of events holds with high probability if the probability of the events tends to unity. We define $\mathcal{S}_{1}$ as the features of the labeled calibration dataset. All functions that we minimize can readily be verified to be continuous, and thus attain a minimum over the compact domains over which we minimize them; thus all our minimizers will be well-defined. We may not mention this further. We denote by $\mathbf{1}[A]$ the indicator of an event $A$. Recall that $\mathcal{H}$ denotes the linear function class =$\mathcal{H}=\{\langle\gamma,\Phi\rangle:\gamma\in\mathbb{R}^{d}\}$. This defines a one-to-one correspondence between $\mathbb{R}^{d}$ and $\mathcal{H}$. This enables us to view functions defined on $\mathbb{R}^{d}$ equivalently as defined on $\mathcal{H}$. In our analysis, we will use such steps without further discussion. Unless stated otherwise, $\mathcal{H}$ is equipped with the norm $\|h\|:=\|\gamma\|_{2}$ for $h=\langle\gamma,\Phi\rangle$. Given a differentiable function $\varphi:\mathcal{H}\to\mathbb{R}$, its directional derivative at $f=\langle\gamma,\Phi\rangle\in\mathcal{H}$ in the direction defined by the function $g\in\mathcal{H}$ is defined as $\frac{d}{d\varepsilon}\big{|}_{\varepsilon=0}\varphi(f+\varepsilon g)$. Note that if we write $g=\langle\tilde{\gamma},\Phi\rangle$ for some $\tilde{\gamma}\in\mathbb{R}^{d}$, then the directional derivative of $\varphi$ at $f$ equals $\langle\tilde{\gamma},\nabla_{\gamma}\varphi(\gamma)\rangle$, where $\nabla_{\gamma}\varphi(\gamma)$ denotes the gradient of $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ evaluated at $\gamma\in\mathbb{R}^{d}$.
Appendix E Conditions
@@
Defining $C_{1}$ as in (10) in Appendix H, assume there exists an upper bound $C_{1,\textnormal{upper}}$ on $\mathbb{E}[C_{1}]$ of constant order.
@@ Appendix F Constants
$A_{5}$ $:=A_{1}+A_{4},\qquad a_{1}:=2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})(1+\beta_{\textnormal{max}}BC_{\Phi}),\qquad a_{2}:=(1-\alpha)C_{\Phi}(C_{1,\textnormal{upper}}+C_{1,\textnormal{max}}).$
@@
$A_{6}:=2\beta_{\textnormal{max}}^{2}\sqrt{4B^{2}\lambda_{\textnormal{max}}(\Sigma)},\qquad A_{7}:=\sqrt{4B^{2}\beta_{\textnormal{max}}^{2}\lambda_{\textnormal{max}}(\Sigma)},\qquad A_{8}:=\sqrt{2B^{2}C_{f}\lambda_{\textnormal{max}}(\Sigma)},\qquad A_{9}:=A_{6}+A_{7},$
@@
$A_{10}$ $:=A_{9}A_{5}^{1/2},\qquad A_{11}:=\max\{A_{9}A_{3}^{1/2},A_{8}A_{5}^{1/2}\},\qquad A_{12}:=A_{9}A_{2}^{1/2},\qquad A_{13}:=A_{8}A_{3}^{1/2},\qquad A_{14}:=A_{8}A_{2}^{1/2}.$
Appendix G Generalization bound for regularized loss
@@
Lemma G.1 (Generalization bound for regularized loss; extension of Shalev-Shwartz and Ben-David, (2014)).
Fix a compact and convex function class $\tilde{\mathcal{H}}$ equipped with a norm $\|\cdot\|_{\tilde{\mathcal{H}}}$, a compact interval $\mathcal{I}\subseteq\mathbb{R}$, and a sample space $\mathcal{Z}$. Consider the objective function $f:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$ given by $(h,\beta,z)\mapsto f(h,\beta,z):=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where $\mathcal{R}:\tilde{\mathcal{H}}\times\mathcal{I}\to\mathbb{R}$ is a regularization function, and $\mathcal{J}:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$ can be decomposed as $\mathcal{J}(h,\beta,z):=\mathcal{J}_{1}(h,\beta,z_{1})+\mathcal{J}_{2}(h,\beta,z_{2})$ for two functions $\mathcal{J}_{1},\mathcal{J}_{2}:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$.
@@
$\mathbb{E}[\mathcal{L}(\hat{h}_{\beta},\beta)-\hat{\mathcal{L}}(\hat{h}_{\beta},\beta)]=(\mathbb{E}[\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1})])+(\mathbb{E}[\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2})]).$
@@
$(\mathbb{E}[(\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1}))\cdot\mathbf{1}[E]]+\mathbb{E}[(\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1}))\cdot\mathbf{1}[E^{c}]])$ (8)
$+(\mathbb{E}[(\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2}))\cdot\mathbf{1}[E]]+\mathbb{E}[(\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2}))\cdot\mathbf{1}[E^{c}]]).$
@@
^$\hat{\mathcal{L}}(h_{2})-\hat{\mathcal{L}}(h_{1})=\hat{\mathcal{L}}_{I,1}(h_{2})-\hat{\mathcal{L}}_{I,1}(h_{1})+\frac{\mathcal{J}_{1}(h_{2},\beta,Z_{I,1})-\mathcal{J}_{1}(h_{1},\beta,Z_{I,1})}{m_{1}}-\frac{\mathcal{J}_{1}(h_{2},\beta,Z_{1}^{\prime})-\mathcal{J}_{1}(h_{1},\beta,Z_{1}^{\prime})}{m_{1}}.$
@@
Appendix H Lipschitz process
Lemma H.1 (Lipschitzness of minimizer of perturbed strongly convex objective).

### v2

@@ 5 Experiments
We compare our method with the following baselines: (1) Split/inductive conformal prediction (Papadopoulos et al., , 2002; Lei et al., , 2018); (2) Weighted-CP: Weighted conformal prediction (Tibshirani et al., , 2019); (3) 2R-CP: The doubly robust method from Yang et al., (2024); (4) DRO-CP: Distributionally robust optimization (Cauchois et al., , 2024); (5) DR-iso: Isotonic distributionally robust optimization (Gui et al., , 2024); (6) Robust-CP: Robust weighted conformal prediction (Ai and Ren, , 2024).
5.1 Choosing the regularization parameter
Equation 4 suggests an optimal choice of the regularization parameter $\lambda$ in the LR-QR algorithm. Guided by this, we form a uniform grid of size ten from $\lambda^{*}/{10}$ to $\lambda^{*}$. We then perform three-fold cross-validation over the combined calibration and unlabeled target datasets (without using any labeled test data) as follows: we train the LR-QR threshold for each $\lambda$, and compute as a validation measure the $\ell^{2}$-norm of the gradient of the LR-QR objective on the held-out fold. We pick $\lambda$ with the smallest average validation measure across all folds.
This validation measure is motivated by our algorithmic development: the first-order conditions of the LR-QR objective play a fundamental role in ensuring valid coverage in the test domain. While the model is trained to satisfy these conditions on the observed data, we seek to ensure this property generalizes well to unseen data. Thus, our selection criterion is based on two key observations: (1) a small gradient of the LR-QR objective implies reliable coverage, and (2) the regularization parameter $\lambda$ balances the generalization error of the two terms in LR-QR. By minimizing this measure, we select a $\lambda$ that optimally trades off these competing factors.
Finally, we re-train the LR-QR threshold on the entire calibration and unlabeled target datasets using this best $\lambda$, and report coverage and interval size on the held-out labeled test set. This ensures that no test labels are used during hyperparameter tuning. Additionally, in Appendix C, we provide deeper insights on different regimes of regularization in practice through an ablation study.
@@ 5.2 Communities and Crime
We evaluate our methods on the Communities and Crime dataset Redmond, (2002), which contains 1994 datapoints corresponding to communities in the United States, with socio-economic and demographic statistics. The task is to predict the (real-valued) per-capita violent crime rate from a 127-dimensional input.
@@
We first randomly select half of the data as a training set, and use it to fit a ridge regression model $\hat{f}$ as our predictor. We tune the ridge regularization with five-fold cross-validation. We use the remaining half to design four covariate shift scenarios, determined by the frequency of a specific racial subgroup (Black, White, Hispanic, and Asian). For each of these features, we find the median value $m$ over the remaining dataset. Datapoints with feature value at most $m$ form our source set, and the rest form our target set. This creates a covariate shift between calibration and test, as the split procedure only observes the covariates and is independent of labels. We then further split the target set into roughly equal unlabeled and labeled subsets. The unlabeled subset and the calibration data (without the labels) is used to estimate $r$, while the labeled test subset is held out only for final evaluation. The same procedure is applied to each of the four racial subgroups, creating four distinct partitions.
@@
Experimental details. The nonconformity score is $s(x,y)\;=\;|y-\hat{f}(x)\,|.$ Several baselines require an estimate of the likelihood ratio $r$, which we obtain by training a logistic regression model $\hat{p}$ to distinguish unlabeled source and target data. We then set $\hat{r}=\frac{\hat{p}}{1-\hat{p}}$, where $\hat{p}(x)$ is the predicted probability that $x$ came from the target distribution. The hypothesis class $\mathcal{H}$ consists of all linear maps from the feature space to $\mathbb{R}$. All experimental results are averaged over 1000 random splits.
@@
Results. Figure 1 displays the results. Notably, split conformal undercovers in two setups and overcovers in the other two. Methods that estimate $r$ and DRO fail to track the nominal coverage, particularly in the first setup on the left. However, the LR-QR method is closer to the nominal level of coverage, showing a stronger adaptivity to the covariate shift.
5.3 Multiple choice questions - MMLU
We evaluate all methods using the MMLU benchmark, which covers 57 subjects spanning a wide range of difficulties. To induce a covariate shift, we partition the dataset by subject difficulty: prompts from subjects labeled as elementary or high school are used for calibration, while those from college and professional subjects form the test set.
Motivated by the design from Kumar et al., (2023), we follow a prompt-based scoring scheme adapted for LLMs: we append the string “The answer is the option:” to the end of each MMLU question and feed the resulting prompt into the Llama 13B model without generating any output. We then extract the next-token logits corresponding to the first decoding position (i.e., immediately after the prompt) and consider the logits associated with the characters A, B, C, and D. These four logits are normalized using the softmax function to produce a probability vector over the answer options.
Experimental details. The nonconformity score is $s(x,y)=1-f(x)_{y}$, where $f(x)_{y}$ is the probability assigned to the correct answer. For $\hat{r}$ and $\mathcal{H}$, we compute prompt embeddings as follows. We extract the final hidden layer outputs from GPT-2 Small to obtain 768-dimensional embeddings. We then apply average pooling across all token embeddings in a prompt to obtain a single fixed-length vector representation for each input. We fit a probabilistic classifier $\hat{p}$ using logistic regression on the unlabeled pooled embeddings from the source and target data, and we set $\hat{r}=\frac{\hat{p}}{1-\hat{p}}$. We set $\mathcal{H}$ to be a linear head on top of the representation layer of the pretrained model.
Results. As shown in Table 1, our LR-QR method achieves near-nominal coverage and has the smallest average prediction set size among methods that achieve approximately 90% or higher coverage, demonstrating both validity and efficiency under covariate shift.
Table 1: Comparison of Methods by Coverage and Set Size
Metric Nominal LR-QR DRO WCP SCP DR-iso Robust-CP 2R-CP
Coverage (%) 90.0 89.6 99.7 86.5 78.1 96.3 95.8 96.9
Set Size 3.38 3.92 3.31 2.60 3.64 3.56 3.80
@@ Appendix A Additional figures
Figure 2: (Above) Coverage, (Below) Average prediction set size.
@@
Appendix B Additional Experiment
B.1 RxRx1 data - WILDS
We consider the RxRx1 dataset Sypetkowski et al., (2023) from the WILDS repository Koh et al., (2021), which is designed to evaluate model robustness under distribution shifts. The RxRx1 task involves classifying cell images based on 1339 laboratory genetic treatments. These images, captured using fluorescent microscopy, originate from 51 independent experiments. Variations in execution and environmental conditions lead to systematic differences across experiments, affecting the distribution of input features (e.g., lighting, cell morphology) while the relationship between inputs and labels remains unchanged. This situation creates covariate shift where the marginal distribution of inputs shifts across domains, but the conditional distribution $\mathbb{P}_{Y|X}$ remains the same.
We use a ResNet50 model He et al., (2016) trained by the WILDS authors on 37 of the 51 experiments. Using the other experiments, we construct 14 distinct evaluations, where each experiment is selected as the target dataset, and its data is evenly split into an unlabeled target set and a labeled test set. The labeled data from the other 13 experiments serves as the source dataset.
Experimental details. The nonconformity score is $s(x,y)=-\log f_{x}(y),$ where $f_{x}(y)$ is the probability assigned the image-label pair $(x,y)$. To estimate $r$, we train a logistic regression model $\hat{p}$ on top of the representation layer of the pretrained model to distinguish unlabeled source and target data, and we set $\hat{r}=\frac{\hat{p}}{1-\hat{p}}$. We set the hypothesis class $\mathcal{H}$ to be a linear head on top of the representation layer of the pretrained model. Experimental results are averaged over 50 random splits.
Results. Figure 4 presents the coverage and average prediction set size for all methods. To enhance visual interpretability, we display results for eight randomly selected settings out of the 14, with the full plot provided in Figure 2. The x-axis shows the indices of the test condition. LR-QR adheres more closely to the nominal coverage value of $0.9$ compared to other methods.
Notably, split conformal prediction, which assumes exchangeability between calibration and test data, shows under- and overcoverage due to the covariate shift. The coverage of weighted CP and 2R-CP is also far from the nominal level, showing that directly estimating the likelihood ratio and conditional quantile is insufficient to correct the coverage violations in the case of high-dimensional image data. Further, the superior coverage of LR-QR is not due to inflated prediction sets.
Appendix C Ablation studies
@@
Appendix D Further related work
@@
Appendix E Notation and conventions
Constants are allowed to depend on dimension only through properties of the population and sample covariance matrices of the features, and the amount of linear independence of the features; see the quantities $\lambda_{\textnormal{min}}(\Sigma)$, $\lambda_{\textnormal{max}}$, $c_{\textnormal{min}}$, $c_{\textnormal{max}}$, and $c_{\textnormal{indep}}$ defined in Appendix F. In the Landau notation ($o$, $O$, $\Theta$), we hide constants. We say that a sequence of events holds with high probability if the probability of the events tends to unity. We define $\mathcal{S}_{1}$ as the features of the labeled calibration dataset. All functions that we minimize can readily be verified to be continuous, and thus attain a minimum over the compact domains over which we minimize them; thus all our minimizers will be well-defined. We may not mention this further. We denote by $\mathbf{1}[A]$ the indicator of an event $A$. Recall that $\mathcal{H}$ denotes the linear hypothesis class =$\mathcal{H}=\{\langle\gamma,\Phi\rangle:\gamma\in\mathbb{R}^{d}\}$. This defines a one-to-one correspondence between $\mathbb{R}^{d}$ and $\mathcal{H}$. This enables us to view functions defined on $\mathbb{R}^{d}$ equivalently as defined on $\mathcal{H}$. In our analysis, we will use such steps without further discussion. Unless stated otherwise, $\mathcal{H}$ is equipped with the norm $\|h\|:=\|\gamma\|_{2}$ for $h=\langle\gamma,\Phi\rangle$. Given a differentiable function $\varphi:\mathcal{H}\to\mathbb{R}$, its directional derivative at $f=\langle\gamma,\Phi\rangle\in\mathcal{H}$ in the direction defined by the function $g\in\mathcal{H}$ is defined as $\frac{d}{d\varepsilon}\big{|}_{\varepsilon=0}\varphi(f+\varepsilon g)$. Note that if we write $g=\langle\tilde{\gamma},\Phi\rangle$ for some $\tilde{\gamma}\in\mathbb{R}^{d}$, then the directional derivative of $\varphi$ at $f$ equals $\langle\tilde{\gamma},\nabla_{\gamma}\varphi(\gamma)\rangle$, where $\nabla_{\gamma}\varphi(\gamma)$ denotes the gradient of $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ evaluated at $\gamma\in\mathbb{R}^{d}$. When it is clear from context, we drop the subscript $\lambda$ from the risks $L_{\lambda}$ and $\hat{L}_{\lambda}$.
Appendix F Conditions
@@
Defining $C_{1}$ as in (8) in Appendix I, assume there exists an upper bound $C_{1,\textnormal{upper}}$ on $\mathbb{E}[C_{1}]$ of constant order.
@@ Appendix G Constants
$A_{5}$ $:=A_{1}+A_{4},\quad a_{1}:=2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})(1+\beta_{\textnormal{max}}BC_{\Phi}),\quad a_{2}:=(1-\alpha)C_{\Phi}(C_{1,\textnormal{upper}}+C_{1,\textnormal{max}}).$
@@
$A_{6}:=2\beta_{\textnormal{max}}^{2}\sqrt{4B^{2}\lambda_{\textnormal{max}}(\Sigma)},\quad A_{7}:=\sqrt{4B^{2}\beta_{\textnormal{max}}^{2}\lambda_{\textnormal{max}}(\Sigma)},\quad A_{8}:=\sqrt{2B^{2}C_{f}\lambda_{\textnormal{max}}(\Sigma)},\quad A_{9}:=A_{6}+A_{7},$
@@
$A_{10}$ $:=A_{9}A_{5}^{1/2},\quad A_{11}:=\max\{A_{9}A_{3}^{1/2},A_{8}A_{5}^{1/2}\},$
$A_{12}$ $:=A_{9}A_{2}^{1/2},\quad A_{13}:=A_{8}A_{3}^{1/2},\quad A_{14}:=A_{8}A_{2}^{1/2}.$
Appendix H Generalization bound for regularized loss
@@
Lemma H.1 (Generalization bound for regularized loss; extension of Shalev-Shwartz and Ben-David, (2014)).
Fix a compact and convex hypothesis class $\tilde{\mathcal{H}}$ equipped with a norm $\|\cdot\|_{\tilde{\mathcal{H}}}$, a compact interval $\mathcal{I}\subseteq\mathbb{R}$, and a sample space $\mathcal{Z}$. Consider the objective function $f:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$ given by $(h,\beta,z)\mapsto f(h,\beta,z):=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where $\mathcal{R}:\tilde{\mathcal{H}}\times\mathcal{I}\to\mathbb{R}$ is a regularization function, and $\mathcal{J}:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$ can be decomposed as $\mathcal{J}(h,\beta,z):=\mathcal{J}_{1}(h,\beta,z_{1})+\mathcal{J}_{2}(h,\beta,z_{2})$ for two functions $\mathcal{J}_{1},\mathcal{J}_{2}:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$.
@@
$\mathbb{E}[\mathcal{L}(\hat{h}_{\beta},\beta)-\hat{\mathcal{L}}(\hat{h}_{\beta},\beta)]=$ $(\mathbb{E}[\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1})])$
$+$ $(\mathbb{E}[\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2})]).$
@@
$(\mathbb{E}[(\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1}))\mathbf{1}[E]]+\mathbb{E}[(\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1}))\mathbf{1}[E^{c}]])$ (6)
$+(\mathbb{E}[(\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2}))\mathbf{1}[E]]+\mathbb{E}[(\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2}))\mathbf{1}[E^{c}]]).$
@@
^$\hat{\mathcal{L}}(h_{2})-\hat{\mathcal{L}}(h_{1})=\hat{\mathcal{L}}_{I,1}(h_{2})-\hat{\mathcal{L}}_{I,1}(h_{1})$ $+\frac{\mathcal{J}_{1}(h_{2},\beta,Z_{I,1})-\mathcal{J}_{1}(h_{1},\beta,Z_{I,1})}{m_{1}}$
$-\frac{\mathcal{J}_{1}(h_{2},\beta,Z_{1}^{\prime})-\mathcal{J}_{1}(h_{1},\beta,Z_{1}^{\prime})}{m_{1}}.$
@@
Appendix I Lipschitz process
Lemma I.1 (Lipschitzness of minimizer of perturbed strongly convex objective).